OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (anonymous):

-5(x+1) < 10 OR -5(x+1) > -10 -5x-5 < 10 -5x-5> -10 -5x < 15 -5x > -5 x < -3 x > 1

OpenStudy (anonymous):

it's wrong cause this sentence \[\left| x+1 \right|\] must be negative which is impossible

OpenStudy (anonymous):

Which one? my answer?

OpenStudy (anonymous):

both of them the question is wrong so don't have any answer

OpenStudy (anonymous):

the sentece that i've mentioned before always could be positive or zero

OpenStudy (anonymous):

-5(x+1) = 10 OR -5(-x-1) = 10 -5x-5 = 10 5x+5 = 10 -5x = 15 5x = 5 x = -3 x = 1

OpenStudy (anonymous):

hey it's not (x+1) it's \[ \left| x+1 \right|\]

OpenStudy (anonymous):

I know , I changed x+1 and -x-1

OpenStudy (anonymous):

no .you can't

OpenStudy (anonymous):

tell me how an absolte value could be negative?

OpenStudy (anonymous):

That's how to solve the inequalities it could be negative or positive

OpenStudy (anonymous):

after you get absolute value from a valuse or sentence the out value is always positive

OpenStudy (anonymous):

\[\left| -2 \right|=2\] \[\left| 0 \right|=0\] \[\left| 2 \right|=2\]

OpenStudy (anonymous):

Can you review the book and how to solve the inequallities?

OpenStudy (anonymous):

for sure the book has example for that

OpenStudy (anonymous):

it's not inequallity

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

it's just a simple equation that include absolute value

OpenStudy (anonymous):

I don't think so but you can tell me how to solve it

OpenStudy (anonymous):

if you see the question there's no inequalitty in the equation

OpenStudy (anonymous):

do you know what's an absolute value?

OpenStudy (anonymous):

Really?! Yes I do for sure

OpenStudy (anonymous):

so what's your problem?

OpenStudy (anonymous):

I don't have problems. it's your post not mine?! Why you posted it and you know how to solve it?

OpenStudy (anonymous):

you can't solve it .in that equation absolute value should be negative to be solved but it's impossible

OpenStudy (anonymous):

ok!!

OpenStudy (anonymous):

!!